# Enhancing multimodal analogical reasoning with Logic Augmented Generation

### v1

@@ Abstract
Recent advances in Large Language Models have demonstrated their capabilities across a variety of tasks. However, automatically extracting implicit knowledge from natural language remains a significant challenge, as machines lack active experience with the physical world. Given this scenario, semantic knowledge graphs can serve as conceptual spaces that guide the automated text generation reasoning process to achieve more efficient and explainable results. In this paper, we apply a logic-augmented generation (LAG) framework that leverages the explicit representation of a text through a semantic knowledge graph and applies it in combination with prompt heuristics to elicit implicit analogical connections. This method generates extended knowledge graph triples representing implicit meaning, enabling systems to reason on unlabeled multimodal data regardless of the domain. We validate our work through three metaphor detection and understanding tasks across four datasets, as they require deep analogical reasoning capabilities. The results show that this integrated approach surpasses current baselines, performs better than humans in understanding visual metaphors, and enables more explainable reasoning processes, though still has inherent limitations in metaphor understanding, especially for domain-specific metaphors. Furthermore, we propose a thorough error analysis, discussing issues with metaphorical annotations and current evaluation methods.
@@ I Introduction
From musical improvisation to solving complex scientific problems, much of what we know and do relies on implicit knowledge: the vast, unconscious reservoir of information we acquire through experience rather than formal instruction [^1] Implicit and tacit knowledge are usually treated as synonyms, for a more comprehensive account on the distinction between the two see [1].. This type of knowledge is employed in analogical reasoning, when we rely on stored patterns, schemas, and relational structures to retrieve and apply knowledge that we may not have explicitly articulated. Therefore, implicit knowledge shapes how we interpret the world, draw inferences, identify conceptual links, and even understand figurative language, all without deliberate effort [2]. Unlike humans, however, machines struggle with these tasks, as they lack direct experience of the physical world, making the computational modeling of implicit learning and consequently analogical reasoning a significant challenge [^2] As shown by the emergence of dedicated conference workshops on the matter, such as the Workshop on Analogical Abstraction in Cognition, Perception, and Language (Analogy-Angle): https://analogy-angle.github.io/.
Figure 1: The LAG framework. The input can be a text or an image, then converted into a FRED-like KG. The graph is then extended with LLM-generated triples according to specific heuristics and defined tasks.
Recent work considers metaphor the highest level in the analogical taxonomy and presents the greatest challenge for Large Language Models (LLMs) [3]. This difficulty arises from the absence of effective prompting techniques to capture the inherent compositionality of language, as well as the lack of anchoring to commonsense knowledge and visual information [4]. Instead of surface-level similarities, analogical reasoning requires the understanding of relational links between concepts, a skill usually learned through the physical interaction with the world [5, 6]. The lack of understanding of relational links in LLMs is a reasoning deficit that remains undetected by standard benchmarks [7], particularly with respect to metaphors. This issue also arises from the lack of a shared problem definition for metaphors and a scarcity of robust, heterogeneous data and evaluation strategies [8].
Although LLMs are trained on diverse corpora that may contain fragments of implicit meaning embedded in human language, analogical reasoning on implicit knowledge is not explicitly codified during training, but might still be represented within LLMs as probabilistic associations across continuous vector spaces. However, relying solely on the probability of association of words to understand and generate language restricts current LLMs [9].
Therefore, despite the multiple remarks that AI systems should be enriched to better mirror deeper semantic and cognitive models of the world, rather than merely relying on correlations between words [10, 11, 12, 13], analogical reasoning remains a significant struggle for the field of Natural Language Processing [14, 13].
@@
Conceptual Blending Theory (CBT) models meaning construction through partially compositional processes in analogy, metaphor, counterfactuals, and related phenomena [15, 16]. Despite its computational promise, studies on CBT largely overlook metaphorical aspects [17, 18, 19] because systematically encoding the vast, unbounded encyclopedic knowledge required for metaphors is challenging [8]. Furthermore, the absence of a core blending ontology has so far limited the computational understanding of these cognitive processes that are, for the most part, innate and implied. Indeed, research shows that LLMs’ performance improves significantly with structured background material [20, 21]. Consequently, combining reliable but labor-intensive semantic technologies with faster, context-dependent yet error-prone LLMs—as demonstrated by Logic Augmented Generation (LAG) in the work by Gangemi and Nuzzolese [22]—opens a promising new avenue for metaphor understanding and analogical reasoning.
@@
This paper proposes a framework based on LAG and inspired by CBT to extract extended Knowledge Graph (KG) triples containing implicit meaning from unlabelled metaphorical data in text and image format (Fig. 1). Integrating structured, machine-interpretable semantics through KGs with fast and dynamic LLM systems allows the real-time flexible understanding and representation of the implied aspects of metaphor in data across various domains and modalities such as news, visual advertisements, memes, and medical sciences.
@@
The contributions of this work are: (i) the adaptation of the LAG framework to enhance multimodal analogical reasoning; (ii) its evaluation on variations on the task of metaphor detection and understanding, namely: linguistic metaphor detection, conceptual metaphor detection and understanding, and visual metaphor understanding; and (iii) a discussion on its computational implementation and applications [^3] All supplementary material, code and data are available at https://github.com/dersuchendee/knowledge-augmented-generation.
@@
The structure of the paper is as follows: Section I introduces the significance and contributions of this work; Section II provides an overview of the theoretical frameworks of analogical abstraction in metaphors; in Section III, we introduce the Blending Ontology for metaphorical conceptual blending. Section IV describes the LAG methodology employed in the paper, Section V and VI concerns its application to three tasks. Section VII analyzes the results, and Section VIII summarizes the contributions and implications for extracting implicit meaning to reason analogically from figurative unlabelled data.
II Background
@@
II-A Implicit knowledge and analogical abstraction
Implicit learning is a process by which knowledge is acquired independently of conscious attempts to do so [23].
Analogical abstraction is a kind of implicitly learned knowledge where as the learner detects and reproduces patterns or analogies across experiences or concepts. This recognition of structural similarities in the process of analogical reasoning can be used to solve problems and make accurate decisions about novel circumstances, and is particularly clear when the connections between concepts are not obvious or are learned through repeated exposure.
@@ II-A Implicit knowledge and analogical abstraction
II-B Conceptual Blending Theory, Metaphor, and Frame Semantics
@@
According to Conceptual Metaphor Theory [26], metaphors create parallels between domains, highlighting experiential correlations. This claim, along with with Mental Spaces Theory, is expanded by Conceptual Blending Theory (CBT), which describes one of the basic cognitive operations of human beings: conceptual combination. According to CBT, in metaphor, a generic space makes abstract and links the two source and target input spaces and generates a new blended space according to a metaphorical criterion or “property” [15]. Indeed, many studies [27, 28, 29] claim that by explaining one object in terms of another, the ‘target’ domain receives a property from the other object, the ‘source’, or in general that the source and target are associated in thought and language because they share a property of some kind. For example, in the metaphor ideas are food, ideas and food are connected through the criterion of internalization (food is inserted and processed in our bodies like ideas), and the two spaces are abstracted with respect to the agent, the input, the process, the output, selectively basing on the context. In this way, the connecting property (internalization) is a higher-level concept that transcends the inherent characteristics of each input space and motivates the comparison.
@@
However, the existence of a shared property is by no means a sufficient condition for two concepts to participate in a metaphorical mapping [30]: for metaphorical meaning to bridge between individuals and within a community, those individuals must already share some understanding of the components involved in the metaphor [31] due to commonsense knowledge and/or specific cultural or social context in which a metaphor is used [32]. In this regard, the CBT framework can be explicitly connected to Fillmore’s frame semantics [33], where certain verbs automatically invoke particular roles, values, and perspectives of discourse participants [34]. Here, different communicative contexts activate background knowledge (i.e. frames) that constrains interpretation. For instance, as shown in Fig. 2, when the conceptual metaphor ideas are food is invoked in the sentence “This is what happens when science serves up facts more indigestible than evolution”, people must already have a background to draw from, some individual grasp of both scientific knowledge and food to understand that it means a way science proposes an idea that is hard to accept from the speaker’s point of view, and that it is a negative outlook on the two domains. The focus-refocus ability of the metaphor is therefore strongly connected to these theoretical frameworks.
Figure 2: The metaphor ideas are food as shown in an example sentence present in the Framester knowledge base [35]. Here, not all roles that exist for each frame mutually map (those that map are in bold and connected through a dotted line), only some are activated in the utterance. Active roles are then subsumed by the roles of the Blending frame (the property), internalization. Overall, the information about the frames is part of the background of the utterance, while the attitude through which we see it is negative.
II-C Theory-integrated metaphor detection and understanding
Recent advancements in metaphor detection have increasingly incorporated theoretical frameworks into computational approaches. Studies leveraging Conceptual Metaphor Theory (CMT) have shown improved baseline results in metaphor identification [36, 37]. For metaphor interpretation, various methods have been developed to discern the source domain implied by the metaphor, given a target domain. These include unsupervised techniques [38], deep learning-based approaches [39], and methods using Large Language Models [40]. However, these methods generally depend on annotations of a single domain rather than exploiting unlabeled data. In Hicke et al., [8], when unlabeled data was exploited, the Metaphor Identification Procedure was preferred as a theoretical framework, thus the metaphors were annotated by category but not by source and target domain. Indeed, the authors recognize this limitation, stating that with the Metaphor Identification Procedure, forcing the models to annotate word-by-word makes it challenging for them to identify metaphors comprising multi-word units.
Other research focuses on extracting properties linking the source and target domains to interpret metaphors [41, 42, 43]. While these approaches yield promising results, they often do not adequately capture the underlying theoretical reasoning processes that are based on the definition of their structural similarity in terms of abstraction and relational dependency on the two metaphorical domains. Additionally, for these systems to work, they require a specified target word.
In the context of visual metaphor understanding, while the work by He et al. [44] focuses on detection, it does not adequately address metaphor understanding according to existing theoretical frameworks, according to which a comprehensive approach should at least involve identifying both source and target domains. To foster richer research in visual metaphors, the Metaclue task set was introduced [45], though its code and dataset are not publicly available. Metaphoric compositions linked by an annotated attribute are available in the ELCo Dataset proposed by Yang et al. [4], where metaphorical composition in emojis is evaluated via a textual entailment task. Integrating heuristics based on existing ontologies on metaphor and knowledge graphs into the metaphor detection and understanding pipeline remains largely unexplored, despite its potential to render theories standardized and machine-readable across multimodal contexts, reduce hallucinations, adapt them for other automated reasoning tasks, and enhance system explainability.
II-D Logic Augmented Generation
@@
The so far described limitations of theory-integrated metaphor detection and understanding studies can be addressed with the integration of ontologies and commonsense KGs in an LLM-based metaphor detection pipeline. The Logic Augmented Generation (LAG) approach, as shown in Gangemi and Nuzzolese [22], conceptualises LLMs as potential Reactive Continuous Knowledge Graphs (RCKGs), which can dynamically adapt to diverse inputs by extending and contextualising Semantic Knowledge graphs (SKGs) that work as base models. LAG exploits the duality between implicit knowledge generated during training and explicit knowledge, externalised in the form of structured knowledge graphs through generation and prompting. RCKG extraction involves a three-step process, i.e. (i) mapping multimodal signal to natural language, (ii) converting natural language into an SKG, and (iii) extending the SKG with tacit knowledge based on multiple heuristics (the extended knowledge graph as XKG).
@@
III The Conceptual Blending Ontology
The Conceptual Blending ontology [^4] Available at http://www.ontologydesignpatterns.org/ont/mdns/blending.owl serves as a heuristic frame of reference in the LAG process. It is based upon mDnS (mereological Descriptions and Situations) [^5] http://www.ontologydesignpatterns.org/ont/mdns/mdns.owl, the Cognitive Perspectivisation ontology as defined in Gangemi and Presutti  [46] [^6] http://www.ontologydesignpatterns.org/ont/persp/perspectivisation.owl, and the MetaNet schema from the Amnestic Forgery ontology by Gangemi et al.  [47].
III-A Background
mDnS is a novel axiomatization for the D&S ontology pattern [^7] http://www.ontologydesignpatterns.org/cp/owl/descriptionandsituation.owl, fully compatible with classic D&S and constructive D&S (cDnS) and is based on intensional/extensional mereology and optimizing expressivity with respect to low complexity for a hybrid reasoner. By introducing the concepts of descriptions and situations, the Description and Situations ontology design pattern  [48] provides a formalization of frame semantics where a situation is a set of facts as described by an observer that is interpreted by a description, a schema that defines the entities observed in a situation.

### v2

@@ Abstract
Recent advances in Large Language Models (LLMs) have demonstrated their capabilities across a variety of tasks. However, automatically extracting implicit knowledge from natural language remains a significant challenge, as machines lack direct experience with the physical world. Given this scenario, semantic knowledge graphs can serve as conceptual spaces that guide LLMs to achieve more efficient and explainable results. In this paper, we apply a Logic Augmented Generation (LAG) framework that leverages the explicit representation of a text through a semantic knowledge graph and applies it in combination with prompt heuristics to elicit implicit analogical connections. This method generates extended knowledge graph triples representing implicit meaning and enabling systems to reason on unlabeled multimodal data regardless of the domain. We validate our work through three metaphor detection and understanding tasks across four datasets, as they require deep analogical reasoning capabilities. The results show that this integrated approach surpasses current baselines and performs better than humans in understanding visual metaphors. It also enables to obtain justifications for the reasoning processes, though still has inherent limitations in metaphor understanding, especially for domain-specific metaphors. Furthermore, we propose a thorough error analysis, discussing issues with existing metaphor datasets, metaphorical annotations, and current evaluation methods.
@@ I Introduction
From musical improvisation to solving complex scientific problems, much of what we know and do relies on implicit knowledge: the vast, unconscious reservoir of information we acquire through experience rather than formal instruction [^1] Implicit and tacit knowledge are usually treated as synonyms. For a more comprehensive account on the distinction between the two see the work by Davies [1].. Consequently, as philosopher Polanyi claimed, “we know more than we can tell” [2]. Analogical reasoning operates on this implicit knowledge base by identifying structural similarities between different domains and shaping how we interpret the world, draw inferences, identify conceptual links, and even understand figurative language [3]. Unlike humans, however, machines struggle with these tasks, as they lack direct experience of the physical world, making the computational modeling of analogical reasoning a significant challenge [^2] As shown by the emergence of dedicated conference workshops on the matter, such as the Workshop on Analogical Abstraction in Cognition, Perception, and Language (Analogy-Angle): https://analogy-angle.github.io/.
Figure 1: Evaluation framework combining automated metrics, expert assessment, and human validation across four layers of analogical reasoning: detection, comprehension, reasoning, and generalization.
Metaphor is arguably the most demanding form of analogical reasoning because it requires not only structural mapping but also the creative blending of distinct conceptual domains [4]. Large Language Models (LLMs) struggle with this task: current prompting methods capture surface‐level word statistics but miss the compositional and experiential grounding needed for genuine analogy [5, 6, 7]. Although these models are trained on vast, heterogeneous corpora that contain fragments of implicit meaning, analogical knowledge is not explicitly codified during training and therefore persists only as probabilistic associations in their embedding space—a dependence that ultimately constrains their reasoning [8]. Their limited grasp of relational links is a known reasoning deficit that standard benchmarks rarely expose [9]. In fact, recent narrative-consistency work shows that, despite stellar token-level scores, frontier LLMs collapse under long-range coherence checks; metaphor requires the same variety of global relational reasoning [10]. Therefore, despite repeated calls to enrich AI systems with deeper semantic and cognitive representations of the world, rather than mere word-level correlations [11, 12, 13, 14], analogical reasoning—especially in the form of metaphor—remains a major open challenge for Natural Language Processing [15, 14]. Progress is further hampered by the absence of a shared problem definition for metaphor and by the scarcity of diverse, well-curated datasets and evaluation protocols [16].
@@
Yet these challenges may be addressable through hybrid neurosymbolic approaches. Research shows that LLMs’ performance improves significantly with structured background material [17, 18]. Consequently, combining reliable but labor-intensive semantic technologies with faster, context-dependent yet error-prone LLMs—as demonstrated by Logic Augmented Generation (LAG) in the work by Gangemi and Nuzzolese [19]—opens a promising new avenue for metaphor understanding and analogical reasoning.
@@
Building on Conceptual Blending Theory—an extension of Lakoff and Johnson’s Conceptual Metaphor Theory that models how multiple mental spaces fuse into novel meaning [20]—we inject an explicit blending ontology into a LAG pipeline (Section IV) to capture the semantic implicit relationships underlying metaphors from unlabelled multimodal data. Integrating structured, machine-interpretable semantics through KGs with fast and dynamic LLMs systems allows the real-time flexible understanding and representation of the implied aspects of metaphor in data across various domains and modalities such as news, visual advertisements, memes, and medical sciences. This hybrid architecture can be classified as a Type 2–3 neurosymbolic system in Kautz’ taxonomy of neurosymbolic systems [21].
@@
The contributions of this work are: (i) the adaptation of the LAG framework to enhance multimodal analogical reasoning; (ii) its multi-dimensional evaluation, as shown in Figure 1, on variations on the task of metaphor detection and understanding, to which we include a novel dataset to test scientific metaphors; and (iii) a discussion on its computational implementation and applications [^3] All supplementary material, code and data are available on GitHub at https://github.com/dersuchendee/metaphorical-logic-augmented-generation. Data is also archived on Zenodo with the following DOI: 10.5281/zenodo.15355852. Our findings show that while we achieve high performance in metaphor detection and generally outperform baselines, the inherent limitations of current LLMs unveil at deeper understanding decreases at higher reasoning levels, with particular challenges in domain-specific and visual metaphors.
@@
The structure of the paper is as follows: Section I introduces the significance and contributions of this work; Sections II and III provide the related works and an overview of the theoretical frameworks of analogical abstraction in metaphors; Section IV describes the LAG methodology employed in the paper, Sections V and VI concern its experimental application to metaphor detection tasks. Section VII discusses the results, and Section VIII summarizes the contributions and implications for extracting implicit meaning to reason analogically from figurative unlabelled data.
II Related works
This section reviews prior work on conceptual blending, metaphor knowledge representation and automated metaphor processing, with a particular focus on detection and interpretation rather than generation.
II-A Foundational approaches to Computational Blending Theory
Early computational approaches to CBT established theoretical foundations that inform current work. Goguen’s algebraic semiotics provided one of the first rigorous mathematical models of conceptual blending, representing metaphors as combinations of structured sign systems and formally specifying how conceptual domains can be merged [22]. Building on this foundation, Kutz and colleagues extended CBT into ontological frameworks, developing methods for algorithmically merging distinct domain ontologies to create novel hybrid concepts through shared structural constraints [23, 24]. Veale’s computational models pioneered practical implementations of CBT, extending metaphor interpretation systems to implement conceptual integration with a focus on generative creativity [25, 26]. While these approaches established crucial theoretical groundwork, they primarily focus on creative concept invention and generation rather than metaphor detection and understanding, with the latter being our primary concern.
II-B Knowledge representation of metaphor
Early knowledge representation efforts handled metaphors in isolation, using SWRL rules to derive new facts from fixed triples [27, 28]. Newer lines of work align metaphors with frame resources: Panayiotou [29] links cross-domain figures through bibliographic ontologies; MetaNet [^4] https://metaphor.icsi.berkeley.edu anchors conceptual metaphors to FrameNet frames; and Amnestic Forgery embeds MetaNet in Framester, yielding 642 mapped metaphors across 270 source and 221 target frames [30]. Despite this progress, static taxonomies are inherently incomplete and need to be complemented by the intervention of systems that can process metaphors dynamically.
II-C Theory-integrated metaphor detection and understanding
Recent advancements in metaphor detection have increasingly incorporated theoretical frameworks into computational approaches. Studies leveraging Conceptual Metaphor Theory (CMT) have shown improved baseline results in metaphor identification [31, 32]. For metaphor interpretation, various methods have been developed to discern the source domain implied by the metaphor, given a target domain. These include unsupervised techniques [33], deep learning-based approaches [34], and methods using Large Language Models [35]. However, these methods generally depend on annotations of a single domain rather than exploiting unlabeled data. In Hicke et al., [16], when unlabeled data was exploited, the Metaphor Identification Procedure was preferred as a theoretical framework, thus the metaphors were annotated by category but not by source and target domain. Indeed, the authors recognize this limitation, stating that with the Metaphor Identification Procedure, forcing the models to annotate word-by-word makes it challenging for them to identify metaphors comprising multi-word units. Other research focuses on extracting properties linking the source and target domains to interpret metaphors [36, 37, 38]. While these approaches yield promising results, they often do not adequately capture the underlying theoretical reasoning processes that are based on the definition of their structural similarity in terms of abstraction and relational dependency on the two metaphorical domains. Additionally, for these systems to work, they tend to require a specified target word. In the context of visual metaphor understanding, while the work by He et al. [39] focuses on detection, it does not adequately address metaphor understanding according to existing theoretical frameworks, which claim that a comprehensive approach should at least involve identifying both source and target domains. To foster richer research in visual metaphors, the MetaCLUE task set was introduced [40], though its code and dataset are not publicly available. Metaphoric compositions linked by an annotated attribute are available in the ELCo Dataset proposed by Yang et al. [5], where metaphorical composition in emojis is evaluated via a textual entailment task. Integrating heuristics based on existing ontologies on metaphor and knowledge graphs into the metaphor detection and understanding pipeline remains largely unexplored, despite its potential to render theories standardized and machine-readable across multimodal contexts, reduce hallucinations, adapt them for other automated reasoning tasks, and enhance system explainability.
II-D Hybrid approaches for computational metaphor processing
In recent years, researchers have grafted structured semantic knowledge onto Transformer backbones to improve metaphor detection and interpretation. FrameBERT [41] fine-tunes RoBERTa while jointly learning embeddings for FrameNet lexical units, yielding interpretable token–frame alignments. Lieto et al. introduce METCL—a TCL-based system that builds a single, emergent prototype for each metaphor by composing head and modifier concepts [42]. The approach is anchored in the FrameNet-style frames embedded in MetaNet and therefore inherits the blind spots of that inventory: if no source/target frame exists, METCL must first extend MetaNet with heuristic candidates before it can even run its logic. METCL’s prototypes are assembled from ConceptNet triples, and risk importing properties that are irrelevant to the discourse at hand.
III Theoretical foundations
@@
III-A Implicit knowledge and analogical abstraction
Implicit knowledge is a type of knowledge acquired independently of conscious attempts to do so and that goes beyond the ability to explain it [43, 2].
Analogical abstraction is a kind of implicitly learned knowledge where patterns or analogies are detected across experiences or concepts. This recognition of structural similarities in the process of analogical reasoning can be used to solve problems and make accurate decisions about novel circumstances, and is particularly clear when the connections between concepts are not obvious or are learned through repeated exposure.
@@ III-A Implicit knowledge and analogical abstraction
III-B Conceptual Blending Theory, metaphor, and frame semantics
@@
According to Conceptual Metaphor Theory [46], metaphors create parallels between domains, highlighting experiential correlations. This claim, along with Mental Spaces Theory, is expanded by CBT, which describes one of the basic cognitive operations of human beings: conceptual combination. According to CBT, in metaphor, a generic space makes abstract and links the two source and target input spaces and generates a new blended space according to a metaphorical criterion or “property” [20]. Indeed, many studies [47, 48, 49] claim that by explaining one object in terms of another, the ‘target’ domain receives a property from the other object, the ‘source’, or in general that the source and target are associated in thought and language because they share a property of some kind. For example, in the metaphor ideas are food, ideas and food are connected through the criterion of internalization (food is inserted and processed in our bodies like ideas), and the two spaces are abstracted with respect to the agent, the input, the process, the output, selectively basing on the context. In this way, the connecting property (internalization) is a higher-level concept that transcends the inherent characteristics of each input space and motivates the comparison.
@@
However, the existence of a shared property is by no means a sufficient condition for two concepts to participate in a metaphorical mapping [50]: for metaphorical meaning to bridge between individuals and within a community, those individuals must already share some understanding of the components involved in the metaphor [51] due to commonsense knowledge and/or specific cultural or social context in which a metaphor is used [52]. In this regard, the CBT framework can be explicitly connected to Fillmore’s frame semantics [53], where certain verbs automatically invoke particular roles, values, and perspectives of discourse participants [54]. Here, different communicative contexts activate background knowledge (i.e. frames) that constrains interpretation. This frame-shifting mechanism central to metaphorical interpretation is therefore well-explained by these theoretical frameworks.
III-C Logic Augmented Generation
@@
The so far described limitations of theory-integrated metaphor detection and understanding studies can be addressed with the integration of ontologies and commonsense KGs in an LLM-based metaphor detection pipeline. The LAG approach, as shown in Gangemi and Nuzzolese [19], conceptualises LLMs as potential Reactive Continuous Knowledge Graphs (RCKGs), which can dynamically adapt to diverse inputs by extending and contextualising Semantic Knowledge graphs (SKGs) that work as base models. LAG exploits the duality between implicit knowledge generated during training and explicit knowledge, externalised in the form of structured knowledge graphs through generation and prompting. RCKG extraction involves a three-step process, i.e. (i) mapping multimodal signal to natural language, (ii) converting natural language into an SKG, and (iii) extending the SKG with tacit knowledge based on multiple heuristics (the extended knowledge graph as XKG).
@@
In summary, our LAG approach represents a Type 2-3 neurosymbolic system [21] that dynamically generates domain-specific knowledge graphs and ontology-based heuristics to guide LLM reasoning, rather than relying on static knowledge bases or simple knowledge retrieval like other hybrid approaches. Unlike purely symbolic approaches that lack flexibility in handling novel metaphors, or purely neural approaches that lack explainability in their reasoning processes, LAG bridges this gap by using structured knowledge representations to guide neural generation while maintaining interpretability. The integration of CBT with SKGs enables our framework to handle metaphor processing task across both textual and visual modalities, addressing the limitations of existing approaches that typically focus on single tasks or modalities. Furthermore, by generating explicit knowledge graphs that represent the blending process, our approach provides explanatory capabilities that current LLM-based methods lack, making the analogical reasoning process transparent and verifiable.
IV Methodology
The proposed method, based on LAG [19], aims to enhance the detection and understanding of metaphor by establishing a standardized and explainable model for the semantic representation of implicit natural language. In this way, the tacit know-how knowledge, following Polanyi’s definitions, can become know-that [2]. Our approach has two main components: i) the representation of inputs, regardless of their modality, into natural language, ii) their conversion into a SKG with Text2AMR2FRED and iii) the SKG enhancement driven by CBT applied to metaphor.
Figure 2: Proposed LAG-based pipeline for the sentence “Crime has infected communities everywhere”.
IV-A Automated multimodal representation with Text2AMR2FRED
AMR identifies the cause of the infection in the sentence, which is, however, not a germ, but crime and the thing that gets infected (community) everywhere. In the FRED-like KG, this graph is enriched with knowledge from Framester such as frames, WordNet synsets and PropBank roles, along with alignment to the DOLCE foundational ontology [58]. However, there is no representation of metaphorical knowledge yet at this stage. While a conceptual metaphor KG sourced from Metanet is contained in Framester, it is still necessary to be able to detect figurative language in any type of data. An image can also be used by being automatically described in natural language. In this case, language is an amodal way to bridge modalities in the resulting supramodal SKG. The SKG can then be extended with implicit contextual knowledge with the support of an LLM.
IV-B Logic Augmented Generation with heuristics
As shown in Fig. 2, the result of the framework that generates XKG triples is the extension of the SKG with new classes and properties according to a specific ontology in the prompt functioning as heuristics. In this work, we adopt the Blending Ontology for this purpose.
IV-B1 The Blending Ontology